How much salary does a man recieves in 9th year if he is starting salary was 2000 with an increament of 200 in each year.
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
That means his salaries are-
2000, 2200, 2400, 2600,.......
Hence this is an AP
a = 2000, d = 200
Now
T9 = a + (n-1) d
T9 = 2000 + 8 * 200 = 3600 Rs
 
Regards
Arun